OpenStudy (anonymous):

Please help!, will medal! Can you tell me if i am right? Which fraction is equivalent to 7 5/9? A. 41/9 B. 75/3 C. 75/9 D. 68/9

OpenStudy (anonymous):

I think that it is A

pooja195 (pooja195):

A is wrong turn the fraction into an improper fraction

pooja195 (pooja195):

denominator x unit + numerator

OpenStudy (fortytherapper):

A is equal to 4 5/9

pooja195 (pooja195):

Hence why I said it was incorrect.

OpenStudy (anonymous):

So D?

pooja195 (pooja195):

D is correct

OpenStudy (unklerhaukus):

7 5/9 = 7 + 5/9 = 7*9/9 + 5/9 = (7*9+5) / 9

OpenStudy (anonymous):

Can you help me with one @UnkleRhaukus ?

OpenStudy (unklerhaukus):

one =1?

OpenStudy (anonymous):

i mean a question XD

OpenStudy (unklerhaukus):

Which question?

OpenStudy (anonymous):

Which mixed number is equivalent to 904/100? A. 9 2/5 B. 9 4/10 C. 9 1/25 D. 9 4/25

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

904/100 = (900+4) / 100 = 900/100 + 4/100 =

OpenStudy (anonymous):

So what would the answer be?

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

= 900/100 + 4/100 = (9*100)/(1*100) + (4*1)/(4*25) =

OpenStudy (anonymous):

i got 226/25 :/

OpenStudy (fortytherapper):

May I ask how you got 226/25? Just to see steps

OpenStudy (anonymous):

well i used a calculator because i am in a hurry

OpenStudy (anonymous):

but i could try it myself

OpenStudy (anonymous):

so if i divide the numerator (the 904) by the denominator (the 100). That'll will give me 9, with a remainder of 4.

OpenStudy (anonymous):

i think the mixed number can be created by using the 9 as the whole number, the 4 as the numerator and the 100 as the denominator?

OpenStudy (anonymous):

So, 904/100 as an improper fraction is 9 4/10

OpenStudy (anonymous):

oki doki i did it

OpenStudy (unklerhaukus):

you mean 9 + 4/100

OpenStudy (fortytherapper):

Almost You made a new denominator when it should have been kept at 100

OpenStudy (anonymous):

oops, sorry im rushing it a lot

OpenStudy (fortytherapper):

No problem. Now all you have to do is simplify